what is y=-3x-4

y=-x+2
Define points x: y:

Answer by stanbon(75887)

what is
y=-3x-4
y=-x+2
-----
Substitute for "y" and solve for "x"::
-3x-4 = -x+2
--------
2x = -6
x = -3
--------
Solve for "y":
y = -x+2
y = --3+2 = 5
--------------
Cheers,
Stan H.
==============